OpenStudy (anonymous):

Suppose that a certain strain of pea plant requires 14 days to reach a height of 6 inches and 30 days to reach a height of 16 inches. How many days would it take a plant of this strain to reach a height of 12 inches? * i got 15 days, am i right? *

OpenStudy (anonymous):

Let's just check with the steps, Did you get m ?

OpenStudy (anonymous):

m?

OpenStudy (anonymous):

the slope :)

OpenStudy (anonymous):

ohh, noo i just cross multiplied and divided (:

OpenStudy (anonymous):

It's okay, you must find the slope to find the equation of the line :) m = y2-y2/x2-x1 after that, choose any of the given points to plug in the following equation : y - yo= m(x-xo) :)

OpenStudy (anonymous):

oh wait a min, wasn't the question find the linear equation?.-.

OpenStudy (anonymous):

noo (:

OpenStudy (anonymous):

oh dear, but you wrote that in the previous one >_< lol

OpenStudy (anonymous):

the other one was, but not this one lol

OpenStudy (anonymous):

oh okay, then you have to use the equation you got in the previous question then plug in x to find y :) since we said that y = height x = number of days when you plug in x, you'll get y, that's it ^_^

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

so you have an equation in the form of : y = mx + B <-- from the previous question plug in y = 12, to get x = number of days ^_^

OpenStudy (anonymous):

what what? ._.

OpenStudy (anonymous):

ohh dont worry, I guess that the question forgot to put "assume a linear eqn "

OpenStudy (anonymous):

lol , anyhow, did you understand what you have to do now bailey ? ^_^

OpenStudy (anonymous):

because otherwise you would assume expxonential growth , but im assuming the asker hasnt done that yet :P

OpenStudy (anonymous):

>_> :P

OpenStudy (anonymous):

do you know what to do now bailey? :)

OpenStudy (anonymous):

kind of (:

OpenStudy (anonymous):

:) not convincing me here lol Alright, did you find the equation from the previous question? :)

OpenStudy (anonymous):

from this question i got y=x+12

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

alright, now we said that x = represents the number of days right? and y =height, right?

OpenStudy (anonymous):

yepp!

OpenStudy (anonymous):

alright then, the question wants to know how many days would it take a plant to reach 12 inches right? so in this case y = 12, all you'vegot to do is plug it in the equation and find x^_^

OpenStudy (anonymous):

got it? :)

OpenStudy (anonymous):

so to find x, what would i have to do, combine like trems? or no.

OpenStudy (anonymous):

nope you have : y = x + 12 where x = days and y = height = 12 12 = x + 12 x = 0....hmm there's something wrong, are you sure of your equation?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

(14,6) ( 30,16) m = (16-6) / (30-14) = (5/8)

OpenStudy (anonymous):

m = y2-y1/x2-x1 m = (16-6)/(30 - 14) m = 0.625

OpenStudy (anonymous):

taake the first point

OpenStudy (anonymous):

y-6 = (5/8) ( x-14)

OpenStudy (anonymous):

so it's y=5/8x+12??????????

OpenStudy (anonymous):

y= (5/8)x -(11/4)

OpenStudy (anonymous):

now plug this in the following : y - yo= m(x-xo) <-- choose any point: y - 12 = 0.625(x-30) y = 0.625x - 6.75

OpenStudy (anonymous):

we are both wrong I think

OpenStudy (anonymous):

that is your equation, now plug in the 12 inches to find the number of days : 12 = 0.625x - 6.75 0.625x = 18.75 x = 30 So it takes 30 days? .-. LOL there's something wrong >_<?

OpenStudy (anonymous):

IKR i couldnt find the answer either

OpenStudy (anonymous):

what are we missing o_o?

OpenStudy (anonymous):

wait I am right :P

OpenStudy (anonymous):

IDK (: lol

OpenStudy (anonymous):

y= (5/8)x -(11/4)

OpenStudy (anonymous):

are you sure x = represents the number of days?, oh so my answer is true? LOL

OpenStudy (anonymous):

^ thats right

OpenStudy (anonymous):

same thing, but mine is in decimal lol, but wait 11/4?

OpenStudy (anonymous):

12 = (5/8)x -11/4 (5/8)x = 14.75 x= (8/5) * 14.75 = 23.6 days

OpenStudy (anonymous):

which makes sense, its half way between 14 and 30

OpenStudy (anonymous):

about

OpenStudy (anonymous):

oh right >_< I've put 12 insteadof 16, sorry sorry , thank you elec ^_^"

OpenStudy (anonymous):

I always miss something o_o, my apologies ^_^

OpenStudy (anonymous):

thanks darlings, your AMAZING (: